OpenStudy (i_love_my_nieces):

HELP FAN AND MEDAL! Determine which ordered pair is a solution of y = –5x + 10. (1 point) (–15, 5) (2, 0) (–1, –5) (3, 25)

OpenStudy (i_love_my_nieces):

@tHe_FiZiCx99

OpenStudy (anonymous):

What you would have to do is plug in the x and y values for each ordered pair. So: -5(-15) + 10 -5(2) + 10 -5(-1) + 10 -5(3) +10

OpenStudy (i_love_my_nieces):

85 0 15 -5 @mipie13

OpenStudy (anonymous):

Okay, so which one of those equals its y value?

OpenStudy (i_love_my_nieces):

-5(-1) + 10?

OpenStudy (i_love_my_nieces):

@mipie13

OpenStudy (i_love_my_nieces):

@mipie13

OpenStudy (anonymous):

No, because look: -5(-1) + 10 Does NOT equal -5.

OpenStudy (i_love_my_nieces):

So, D

OpenStudy (anonymous):

Having really bad lag, hold on please.

OpenStudy (i_love_my_nieces):

Same here. OK.

OpenStudy (anonymous):

I'm back. Let me show you: -5(-15) + 10 = 85 Ordered Pair: (-15, 5) Y-Axis is 5 -5(2) + 10 = 0 Ordered Pair: (2, 0) Y-Axis is 0 -5(-1) + 10 = 15 Ordered Pair: (-1, -5) Y-Axis is -5 -5(3) +10 = -5 Ordered Pair: (3, 25) Y-Axis is 25 Remember, we are trying to find which of these completes the equation. Which of the answer choices matches up with its Y-AXIS?

OpenStudy (i_love_my_nieces):

So therefore it's C

OpenStudy (anonymous):

Let me try it a different way, Which of the answer choices has the same number all the way across? A. 85, 5, 5. B. 0, 0, 0. C. 15, -5, -5. D. -5, 25, 25.

OpenStudy (i_love_my_nieces):

A?

OpenStudy (anonymous):

:P... It's B.